OpenStudy (anonymous):

Evaluate if k= -1/2: 8- k^3 ------ k-2

OpenStudy (theviper):

Put the value of k=-1/2 in the question & solve ;)

Parth (parthkohli):

That'd get\[8 - \left({-1 \over 2} \right)^3 \over k -\left( -1 \over 2\right)\]

Parth (parthkohli):

You can also use "difference of cubes".\[a^3 - b^3 \implies (a + b)(a^2 - ab + b^2) \]

OpenStudy (theviper):

\[\frac{8+\frac{1}{8}}{\frac{-1}{2}-2}\]

Parth (parthkohli):

Sorry. I meant\[{-1 \over 2} -2\]in the denominator.

OpenStudy (anonymous):

so is the answer 3.25?